OpenStudy (anonymous):

Simplify and leave your answer as a proper fraction or a mixed number. 2 and 1/2 + 2 and 1/5+3 and 1/10. Enter your answer as a reduced mixed number or as a whole number

OpenStudy (anonymous):

2 1/2 = 5/10 2 1/5 = 2/10 3 1/10 = 1 10 -------------add 7 8/10 = 7 4/5

OpenStudy (anonymous):

so the answer is 7 and 4/5

OpenStudy (anonymous):

yes